# Beneath the Surface: How Large Language Models Reflect Hidden Bias

### v1

@@ Abstract
The exceptional performance of Large Language Models (LLMs) often comes with the unintended propagation of social biases embedded in their training data. While existing benchmarks evaluate overt bias through direct term associations between bias concept terms and demographic terms, LLMs have become increasingly adept at avoiding biased responses, creating an illusion of neutrality. However, biases persist in subtler, contextually hidden forms that traditional benchmarks fail to capture. We introduce the Hidden Bias Benchmark (HBB), a novel dataset designed to assess hidden bias that bias concepts are hidden within naturalistic, subtly framed contexts in real-world scenarios. We analyze six state-of-the-art LLMs, revealing that while models reduce bias in response to overt bias, they continue to reinforce biases in nuanced settings. Data, code, and results are available at https://github.com/JP-25/Hidden-Bias-Benchmark.
Beneath the Surface: How Large Language Models Reflect Hidden Bias
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) is frequently accompanied by the propagation of social bias inherent in their training data Gallegos et al. (2024a); Hofmann et al. (2024); Navigli et al. (2023); Cui et al. (2024). These biases raise serious ethical concerns, as they perpetuate stereotypes, reinforce discrimination, and negatively impact real-world decision-making. In domains such as hiring, law enforcement, and content moderation, the use of these models in real-world applications may disproportionately harm marginalized individuals and communities Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023).
Figure 1: Hidden Bias Benchmark (HBB) reflects bias.
Numerous studies Parrish et al. (2022); Marchiori Manerba et al. (2024); Nangia et al. (2020); Nadeem et al. (2021) benchmark Overt Bias in LLMs by analyzing direct associations between a specific demographic term and a bias-related concept term. As illustrated in Figure 1, example (a) from BBQ Parrish et al. (2022) can demonstrate overt bias when the model consistently associates “Margaret” (female) with the term “bad at math” and “George” (male) with the term “good at math”, or vice versa. However, a fundamental issue remains: overt bias can be simply mitigated by breaking the direct association between demographic terms and concept terms Gallegos et al. (2024b); Li et al. (2024). Additionally, as LLMs evolve, their responses to overt bias evaluations have become more neutral and self-regulated, frequently aligning with socially desirable norms. This trend is largely driven by advances in model training techniques, particularly instruction tuning and alignment strategies, which encourage neutrality in responses to overtly biased contexts Ouyang et al. (2022); Zhang et al. (2023); Peng et al. (2023); Ji et al. (2024). Consequently, existing overt bias benchmarks often report low bias scores for LLMs. In our experiments (details in Section 4.2.2), GPT-4o achieves a score of -0.000807 on the BBQ-ambiguous dataset, with 0 indicating no bias.
In real-world scenarios, biases are hidden within context rather than overtly stated. Typically, associations between demographic terms and bias-related concept terms are concealed within contexts, without explicitly referencing them. Specifically, bias-related concepts are usually reflected through depictions of personality traits, actions, behaviors, emotions, and more. Meanwhile, demographic identities can be subtly conveyed through indirect descriptors. We define this phenomenon as Hidden Bias, where biases are behind the scenes, manifesting through associations between hidden descriptions of demographic identities and concepts within real-world scenarios, without overt reference. As shown in Figure 1 example (b), within the same scenario, the male identity is subtly indicated by the name “George”, while the female identity is represented by “Margaret”. Option A portrays behaviors that implicitly convey the concept of “bad at math”, whereas Option B reflects the notion of “good at math”. Hidden bias arises when females are consistently associated with the concept depiction of “bad at math” while males are linked to the notion of “good at math”, or vice versa.
To bridge this gap, we propose the Hidden Bias Benchmark (HBB), a systematic framework for evaluating hidden bias through structured test instances. Each test instance in HBB consists of a pair of questions, as illustrated in example (b) of Figure 1. As demonstrated, LLMs reinforce stereotypes when biases are subtly hidden within realistic scenarios. For instance, while an LLM may reject a direct stereotype (e.g., Figure 1 (a)), it may still unintentionally perpetuate the same bias when the contexts are reframed in a more subtle, contextually hidden manner (Figure 1 (b)). In our experiments (details in Section 4.2.2), when we use our HBB to examine the same set of biases tested by BBQ, we observe a significant increase in bias metrics for GPT-4o, illustrating the necessary and significance of investigating the proposed hidden bias.
As LLMs become more adept at recognizing and avoiding overt bias, evaluating how models respond to contexts with subtly hidden bias becomes increasingly crucial. Our HBB provides a comprehensive framework for examining biases that persist despite overt bias avoidance mechanisms, offering a more robust evaluation of bias in LLMs. Data, code, and results are available at https://github.com/JP-25/Hidden-Bias-Benchmark. In summary, our contributions are threefold:
@@
    We conceptualize hidden bias in LLMs by focusing on biases that measure the association between hidden demographic descriptors and bias-related concept descriptions.
@@
    Our HBB spans five key social categories: Age (4,641 test instances), Gender (6,188 test instances), Race Ethnicity (Race) (61,880 test instances), Socioeconomic Class (SES) (3,094 test instances), and Religions (27,846 test instances). In addition to the original Multiple-Choice-Question (MCQ) version of HBB, we also introduce a Semi-Generation-based HBB (HBB-SG). HBB-SG is motivated by the increasing application of LLMs in open-ended generative tasks, providing a more realistic assessment of hidden bias in generation settings.
@@
    We evaluate hidden bias that previous works cannot measure across six LLMs, analyzing bias patterns across models, demographic categories, identities, and descriptors to offer a comprehensive view of how LLMs perpetuate hidden bias. Notably, we find that more advanced models, such as GPT-4o, exhibit higher hidden bias while showing lower overt bias.
@@ 2 Related Work
Overt Bias Benchmarks.
Overt bias in LLMs has been widely examined using benchmarks that assess model preference for stereotypical over anti-stereotypical associations when explicit concept terms with demographic identities. And multiple benchmarks have been designed to quantify overt bias from diverse perspectives, facilitating structured evaluations Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023). These benchmarks establish the foundation for overt bias evaluation, assessing how LLMs respond to overtly biased statements.
Hidden Bias and the Evolution of Model Behavior.
As LLMs advance, their responses to overt bias evaluations have become more neutral and self-regulated, often producing answers that align with socially desirable norms. Consequently, traditional overt bias benchmarks mentioned previously, often show reduced bias scores for LLMs. However, biases may persist in subtler, more hidden ways that traditional evaluation methods fail to capture Bai et al. (2024); Smith et al. (2022). Our proposed Hidden Bias Benchmark (HBB) evaluates hidden bias by analyzing response variations across parallel test instances with different demographic descriptors, where biases are subtly hidden in naturalistic language. More extensive discussions of related works are provided in Appendix B.
3 Hidden Bias Benchmark (HBB)
Figure 2: Hidden Bias Benchmark (HBB) workflow.
As LLMs do not show a high level of bias in existing overt bias benchmarks, we aim to develop a dataset specifically designed to measure hidden bias in LLMs that previous works do not measure.
@@ 3.1 Dataset Generation
Figure 2 illustrates the complete workflow for dataset construction. We first extract bias summary with explanations from existing datasets, which are then used to generate a pair of opposite bias-related concepts. Next, we prompt GPT-4o to generate a real-life contextual scenario and two answer options, subtly hiding the pair of opposite bias-related concepts within the sentences. The resulting questions, incorporating various demographic identities, form our proposed Hidden Bias Benchmark (HBB). Additionally, all generated contexts and answer choices undergo manual review to ensure logical consistency and accurate representation of pairs of opposite bias-related concepts. The detailed design of each step is presented in the following sections.
3.1.1 Pairs of Opposite Bias Related Concepts
@@
The identification of bias concepts is fundamental to understanding social bias. For instance, specific occupations are often stereotypically linked to either men or women. We compile these bias concepts from well-established overt social bias datasets, including BBQ Parrish et al. (2022), SOFA Marchiori Manerba et al. (2024), CrowS-Pairs (CS)Nangia et al. (2020), and StereoSet (SS)Nadeem et al. (2021). However, the previous works contain bias concepts directly but often contain stereotypes that are low-quality Blodgett et al. (2021), overly specific, or lack the opposite counterpart, and no works systematically extract all concepts comprehensively. In this regard, we are able to extract high-quality concepts by employing GPT-4o in a structured three-step process:
Bias Summary
@@
In general, as shown in Figure 2, GPT-4o is prompted to process inputs from previously mentioned datasets, such as BBQ, using a given context and question. The bias concept in BBQ is embedded within the question, such as: “Who is good at math?”. As a result, the generated bias summary states: “The question implies a stereotype that one gender is inherently less skilled at math than the other.” The complete prompts for each dataset, with slight variations in input formats, are provided in Table 12 in Appendix C.1.1.
Raw Concept Pairs
@@
Using the bias summary from the previous step, we construct a new prompt for GPT-4o, incorporating a few examples to facilitate in-context learning Brown et al. (2020). This approach allows GPT-4o to identify general bias concepts that reflect traditional biases, paired with their corresponding opposite bias concepts. Consequently, we generate raw concept pairs, each containing a bias concept and the opposite bias concept. Referring to Figure 2, the raw concept pair is “good at math; bad at math”. The full set of prompts is provided in Table 13 in Appendix C.1.2.
Post-hoc Check
@@
Finally, we employ GPT-4o for a final quality check, reviewing the generated concept pairs alongside their corresponding bias summary to ensure logical consistency, relevance, and proper alignment with identified biases. If the generated concepts are of low quality or misaligned with their explanations, GPT-4o automatically revises them to enhance consistency and generates a more suitable concept pair. The complete prompts are shown in Table 14 in Appendix C.1.3.
@@ 3.1.2 Question Design
After acquiring high-quality bias concept pairs, we leverage GPT-4o to generate raw questions for the dataset, each paired with a contextual scenario and two corresponding answer options. The question structure follows a simple two-step process:
Context Design
We first omit demographic information from the context to later assess whether certain concepts trigger biases across different demographic identities. With this approach, GPT-4o functions as a story writer, generating a concise sentence that incorporates [[X]] as the main character to depict a real-world scenario with minimal details, forming the context without unnecessary elements. The generated context functions as the opening sentence, providing a scene description with [[X]]. It later guides GPT-4o in generating a sentence that depicts the bias concept followed by this context. And [[X]] will be replaced with different demographic identities during data construction in Section 3.1.3. As demonstrated in Figure 2, GPT-4o generates a simple and plain context scene without any extra information“[[X]] sat at the dining table, surrounded by textbooks and notes.” The complete prompts for context design are shown in Table 15 in Appendix C.2.
Answer Options Design
Next, we continue to utilize GPT-4o as a story generator to expand the narrative based on the provided context, ensuring that [[X]] is described in alignment with one of the concept pairs. For the remaining concepts, we apply the same approach, providing context and prompting GPT-4o to generate a narrative incorporating [[X]] according to the respective concept. In summary, we craft prompts that subtly describe [[X]], deliberately avoiding explicit references to the bias concept. Specifically, answer options (see Option 1 and Option 2 in Figure 2 with [[X]]) should indirectly characterize [[X]] through attributes such as personality traits, behaviors, emotions, decision-making styles, values, and more. The complete prompts for answer options design are shown in Table 15 in Appendix C.2.
@@ 3.1.3 Data Construction
Furthermore, not only the pairs of opposite bias-related concepts can be hidden by descriptions, but the demographic identities can also be hidden by different types of descriptors. Traditional overt bias benchmarks have not comprehensively examined how different demographic identity descriptors can be expressed in varying degrees of explicitness and implicitness. Instead, they use direct demographic identities, such as “the woman” and “the man”. Our work fills this gap by systematically investigating how demographic descriptors for same identity replacements (explicit way and implicit way) affect bias exhibitions in LLMs. And by structuring demographic descriptors from most implicit to most explicit, we ensure that our dataset captures a broad spectrum of potential bias triggers.
Therefore, at this stage, [[X]] is replaced with various subtle demographic descriptors without direct demographic references, ensuring a comprehensive evaluation of hidden bias across multiple identity types. For example, in the bias category of Age, [[X]] for an older identity may be replaced with “a grandmother living in a nursing home”, while for a younger identity, it may be replaced with “a daughter who is a college freshman”. Terms like “retirement” and “Gen-X” further reinforce age representation without explicitly stating “Old” or “Young.” Similarly, for Race Ethnicity, [[X]] is subtly depicted using names, pet phrases, and culturally significant holidays. Gender is represented through terms such as mother/father or professions like actor/actress. For Socioeconomic Class, descriptions of living conditions are used, and religious identity is expressed through references to religious practices and behaviors. Table 10 provides a systematic summary of subtle identity replacements in Appendix C.3, ranging from implicit to explicit identity descriptors, while Table 4 details the randomly assigned names for [[X]].

### v2

# What’s Not Said Still Hurts: A Description-Based Evaluation Framework for Measuring Social Bias in LLMs

@@ Abstract
Large Language Models (LLMs) often exhibit social biases inherited from their training data. While existing benchmarks evaluate bias by term-based mode through direct term associations between demographic terms and bias terms, LLMs have become increasingly adept at avoiding biased responses, leading to seemingly low levels of bias. However, biases persist in subtler, contextually hidden forms that traditional benchmarks fail to capture. We introduce the Description-based Bias Benchmark (DBB), a novel dataset designed to assess bias at the semantic level that bias concepts are hidden within naturalistic, subtly framed contexts in real-world scenarios rather than superficial terms. We analyze six state-of-the-art LLMs, revealing that while models reduce bias in response at the term level, they continue to reinforce biases in nuanced settings. Data, code, and results are available at https://github.com/JP-25/Description-based-Bias-Benchmark.
What’s Not Said Still Hurts: A Description-Based Evaluation Framework for Measuring Social Bias in LLMs
@@ 1 Introduction
The remarkable performance of Large Language Models (LLMs) is frequently accompanied by the propagation of social bias inherent in their training data Gallegos et al. (2024a); Hofmann et al. (2024); Navigli et al. (2023); Cui et al. (2024). These biases raise serious ethical concerns by perpetuating stereotypes and reinforcing discrimination. In many high-stakes domains, the deployment of LLMs may disproportionately harm marginalized individuals and communities Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023).
Figure 1: Description-based Bias Benchmark (DBB) measures bias that existing term-based benchmarks cannot.
Numerous studies Parrish et al. (2022); Marchiori Manerba et al. (2024); Nangia et al. (2020); Nadeem et al. (2021) evaluate bias from a term-based perspective, which assesses direct lexical-level associations between demographic identities and bias-related terms (demographic-term associations). As illustrated in Figure 1(a), an example from BBQ Parrish et al. (2022) demonstrates bias when the model consistently associates “Margaret” (female) with the term “bad at math” and “George” (male) with the term “good at math”, or vice versa. We call such a bias measurement focused on associations to terms term-based evaluation. However, this evaluation method is too superficial, and the measurement method can be easily manipulated by breaking the direct association between demographic identities and bias-related terms Gallegos et al. (2024b); Li et al. (2024). Thus, state-of-the-art (SOTA) LLMs usually show a low level of bias when evaluated by existing term-based evaluation benchmarks Ouyang et al. (2022); Zhang et al. (2023); Peng et al. (2023); Ji et al. (2024). In our experiments (details in Section 4.2.2), GPT-4o achieves a score of -0.000807 on the BBQ-ambiguous dataset, with 0 indicating no bias. Does this suggest that LLMs are truly unbiased – or, current benchmarks measuring bias in the superficial term-based way are insufficient to capture the full spectrum of biases?
Indeed, real-world biases are rarely expressed in explicit terms. Instead, they operate behind the scenes, manifesting through semantic-level associations between demographic identities and bias-related concepts (demographic-concept associations). Critically, these concepts are conveyed through depictions of behaviors, emotions, mental activities, and more. A single concept can be represented by different descriptions. We call this semantic-level measurement manner description-based evaluation, distinguishing it from the term-based methods in prior work. Within the same scenario as Figure 1(a), Option A in Figure 1(b) portrays behaviors that subtly imply the concept of “bad at math”, whereas Option B reflects the notion of “good at math”. Bias is detected when females are consistently associated with depictions aligned with the concept of “bad at math”, while males are linked to “good at math”, or vice versa.
Existing works on social bias in LLM focused on superficial demographic-term associations and overlooked the evaluation of demographic-concept associations. Toward this, we propose the Description-based Bias Benchmark (DBB) to systematically evaluate social bias. With this new benchmark, we find that LLMs often avoid showing bias at the term level (e.g., Figure 1(a)) but can unintentionally perpetuate the same instance of bias when expressed at the description level (Figure 1(b)). In our experiments (Section 4.2.2), when we use DBB to examine the same set of biases tested by BBQ, we observe a significant increase in bias metrics for GPT-4o, illustrating the necessity and significance of investigating bias in the proposed nuanced setting. Data, code, and results are available at https://github.com/JP-25/Description-based-Bias-Benchmark.
In summary, our contributions are:
@@
    We evaluate social bias in LLMs by focusing on semantic-level associations between demographic identities and bias-related concepts reflected by varying descriptions.
@@
    DBB spans five social categories: Age (4,641 test instances), Gender (6,188), Race Ethnicity (Race) (61,880), Socioeconomic Class (SES) (3,094), and Religions (27,846). Alongside the original Multiple-Choice-Question (MCQ) version, we introduce a Semi-Generation version (DBB-SG). DBB-SG is motivated by the increasing application of LLMs in open-ended generation tasks, providing a more practical assessment of bias in generations.
@@
    We evaluate bias across six LLMs, analyzing bias patterns across models, demographic categories, identities, and descriptors to offer a comprehensive view of how LLMs perpetuate bias in description-based evaluation. Notably, advanced models like GPT-4o exhibit a higher level of bias in the description-based method despite showing a lower level of bias in the term-based approach.
@@ 2 Related Work
Due to space limitations, more extensive discussions of related works are provided in Appendix B.
Bias evaluation in LLMs has been extensively studied via term-based methods that assess lexical-level associations between demographic terms and bias terms (term-based evaluation). Multiple benchmarks Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023); May et al. (2019); Caliskan et al. (2017) quantify bias using term-based evaluation from diverse perspectives and form the foundation for evaluating how LLMs respond to superficially biased content Zhao et al. (2024b); Venkit et al. (2022); Field and Tsvetkov (2020); Lin and Li (2025); Tan and Lee (2025).
As LLMs evolve, their responses to such direct evaluations have become more neutral and socially aligned, often resulting in lower bias scores. However, this trend does not necessarily indicate genuine mitigation. Subtler forms of bias may persist in ways that traditional benchmarks cannot detect Bai et al. (2024); Smith et al. (2022). Bai et al. (2024) move beyond conventional benchmarks by prompting LLMs to associate specific words with demographics, then generate narratives. Nonetheless, the method remains within a term-based mode, relying on lexical-level associations.
In contrast, we introduce the Description-based Bias Benchmark (DBB), which assesses bias by description-based evaluation. This approach captures semantic-level associations between demographic identities and bias-related concepts that are subtly hidden within naturalistic contexts. Specifically, DBB evaluates bias by measuring response variations across parallel test instances with different demographic descriptors where bias-related concepts are hidden within naturalistic contexts rather than superficial term associations.
3 Description-Based Bias Benchmark
Figure 2: Description-based Bias Benchmark (DBB) workflow.
As LLMs show low bias in existing term-based bias benchmarks, we aim to develop a dataset designed to measure bias by the description-based method in LLMs that previous works do not measure.
@@ 3.1 Dataset Generation
Figure 2 illustrates the complete workflow for dataset construction. We first extract a bias summary with explanations from existing datasets, which are then used to generate a pair of opposite bias-related concepts. Next, we prompt GPT-4o to generate a real-life contextual scenario and two answer options, subtly hiding the pair of opposite bias-related concepts within the sentences. The resulting questions, incorporating various demographic identities, form our proposed Description-based Bias Benchmark (DBB). Additionally, all generated contexts and answer choices undergo manual review to ensure logical consistency and accurate representation of pairs of opposite bias-related concepts. The detailed design of each step is presented in the following sections.
3.1.1 Pairs of Opposite Bias-Related Concepts
@@
The identification of bias concepts is fundamental to understanding social bias. For instance, specific occupations are often stereotypically linked to either men or women. We compile these bias concepts from well-established term-based social bias datasets, including BBQ Parrish et al. (2022), SOFA Marchiori Manerba et al. (2024), CrowS-Pairs (CS) Nangia et al. (2020), and StereoSet (SS) Nadeem et al. (2021). However, the previous works contain bias concepts directly but often contain stereotypes that are low-quality Blodgett et al. (2021), overly specific, or lack the opposite counterpart, and no works systematically extract all concepts comprehensively. In this regard, we are able to extract high-quality concepts by employing GPT-4o in a structured three-step process:
@@
Bias Summary. In general, as shown in Figure 2, GPT-4o is prompted to process inputs from previously mentioned datasets, such as BBQ, using a given context and question. The bias concept in BBQ is embedded within the question, such as: “Who is good at math?”. As a result, the generated bias summary states: “The question implies a stereotype that one gender is inherently less skilled at math than the other.” The complete prompts for each dataset, with slight variations in input formats, are provided in Table 12 in Appendix C.1.1.
@@
Raw Concept Pairs. Using the bias summary from the previous step, we construct a new prompt for GPT-4o, incorporating a few examples to facilitate in-context learning Brown et al. (2020). This approach allows GPT-4o to identify general bias concepts that reflect traditional biases, paired with their corresponding opposite bias concepts. Consequently, we generate raw concept pairs, each containing a bias concept and the opposite bias concept. Referring to Figure 2, the raw concept pair is “good at math; bad at math”. The full set of prompts is provided in Table 13 in Appendix C.1.2.
@@
Post-hoc Check. Finally, we employ GPT-4o for a final quality check, reviewing the generated concept pairs alongside their corresponding bias summary to ensure logical consistency, relevance, and proper alignment with identified biases. If the generated concepts are of low quality or misaligned with their explanations, GPT-4o automatically revises them to enhance consistency and generate a more suitable concept pair. The complete prompts are shown in Table 14 in Appendix C.1.3.
@@ 3.1.2 Question Design
After acquiring high-quality bias concept pairs, we leverage GPT-4o to generate raw questions for the dataset, each paired with a contextual scenario and two corresponding answer options. The question structure follows a simple three-step process:
Context Design. We first omit demographic information from the context to later assess whether certain concepts trigger biases across different demographic identities. With this approach, GPT-4o functions as a story writer, generating a concise sentence that incorporates [[X]] as the main character to depict a real-world scenario with minimal details, forming the context without unnecessary elements. The generated context functions as the opening sentence, providing a scene description with [[X]]. It later guides GPT-4o in generating a sentence that depicts the bias concept followed by this context. And [[X]] will be replaced with different demographic identities during data construction in Section 3.1.3. As demonstrated in Figure 2, GPT-4o generates a simple and plain context scene without any extra information“[[X]] sat at the dining table, surrounded by textbooks and notes.” The complete prompts for context design are shown in Table 15 in Appendix C.2.
Answer Options Design. Next, we continue to utilize GPT-4o as a story generator to expand the narrative based on the provided context, ensuring that [[X]] is described in alignment with one of the concept pairs. For the remaining concepts, we apply the same approach, providing context and prompting GPT-4o to generate a narrative incorporating [[X]] according to the respective concept. In summary, we craft prompts that subtly describe [[X]], deliberately avoiding explicit references to the bias concept. Specifically, answer options (see Option 1 and Option 2 in Figure 2 with [[X]]) should indirectly characterize [[X]] through attributes such as personality traits, behaviors, emotions, decision-making styles, values, and more. The complete prompts for answer options design are shown in Table 15 in Appendix C.2.
@@
Manual Quality Evaluation. To ensure the quality of the generated raw data, we manually evaluate 100 randomly sampled raw instances. Each instance is assessed along four dimensions: (1) contextual fluency: the context is grammatically correct and free of awkward phrasing; (2) context-option coherence: both options are logically consistent with the given context; (3) linguistic naturalness: the language in both context and options reads naturally, resembling real-word usage; and (4) semantic alignment: the options reflect the intended bias-related concepts in a hidden descriptive manner rather than through superficially direct expressions.
@@ 3.1.3 Data Construction
Furthermore, not only the pairs of opposite bias-related concepts can be hidden by descriptions, but demographic identities can also be hidden by different types of descriptors. Traditional term-based bias benchmarks have not comprehensively examined how different demographic identity descriptors can be expressed in varying degrees of explicitness and implicitness. Instead, they use direct demographic identities, such as “the woman” and “the man”. Our work fills this gap by systematically investigating how demographic descriptors for same identity replacements (explicit way and implicit way) affect bias exhibitions in LLMs. And by structuring demographic descriptors from most implicit to most explicit, we ensure that our dataset captures a broad spectrum of potential bias triggers.
Thus, at this stage, [[X]] is replaced with various subtle demographic descriptors without direct demographic references, ensuring a comprehensive evaluation of bias across multiple identity types. For example, in the bias category of Age, [[X]] for an older identity may be replaced with “a grandmother living in a nursing home”, while for a younger identity, it may be replaced with “a daughter who is a college freshman”. Terms like “retirement” and “Gen-X” further reinforce age representation without explicitly stating “Old” or “Young.” Similarly, for Race Ethnicity, [[X]] is subtly depicted using names, pet phrases, and culturally significant holidays. Gender is represented through terms such as mother/father or professions like actor/actress. For Socioeconomic Class, descriptions of living conditions are used, and religious identity is expressed through references to religious practices and behaviors. All descriptors are drawn from and inspired by prior works, including BBQ Parrish et al. (2022), SOFA Marchiori Manerba et al. (2024), CS Nangia et al. (2020), and SS Nadeem et al. (2021). Table 10 provides a systematic summary of subtle identity replacements in Appendix C.3, ranging from implicit to explicit identity descriptors, while Table 4 details the randomly assigned names for [[X]].